Erin and Sam are selling fruit far a school fundraiser. Customers can buy boxes of tangerines and boxes of pears. Erin sold 13 boxes of tangerines and 14 boxes of pears for a total of $276. Sam sold 14 boxes of tangerines and 2 boxes of pears for a total of $88. What is the cost each of the one box of tangerines and one box of pears? Show your work.
Answer by addingup(3677)

13t+14p = 276 (1)
14t+2p = 88 (2)
--------------------------------
Multiply equation (2) times 7 and subtract from (1):
13t+14p = 276
-
98t+14p = 616
--------------
-85t = -340
Divide both sides by -85 and remember -/- = +
t = 4 The tangerines are $4 and the pears:
14(4)+2p = 88
56+2p = 88
2p = 32
p = 16 Pears are $16
Let's try our answers in the other equation to make sure we have the correct answer:
13(4)+14(16) = 276
52+224 = 276 Add on left to see that we hare the correct answers.